1. Rank substituents in order of their priority when assigning the E or Z label to an alkene.

CH2CH2Cl

H

I

CH2OH

NH2

2. The following is a multi-step reaction. The rate-limiting step is unimolecular, with A as the sole reactant.

A +B --> C+D

If [A] and [B] are both 0.130 M, then the rate of reaction is 0.0030 M/s.

(a) What is the rate of the reaction if [A] is doubled?

(b) Starting with the original concentrations, what is the rate of the reaction if [B] is doubled?

(c) Starting with the original concentrations, what is the rate of the reaction if [A] and [B] are both increased by a factor of five?

3.

4.
Answer #1

1. Priority: I > NH2 > CH2OH > CH2CH2Cl > H

Rule: 1. Higher priority by atomic number

Rule   2. If fisrt connected atoms to the double bond are same, then check for the next connected atoms like in CH2OH > CH2CH2Cl. Here, the second connected atom is 'O' in CH2OH and 'C' in CH2CH2Cl . Like wise it can be extended to next next next..... atoms untill there is a difference. This rule is known as "first point of difference" rule.

2. Ans.

             The rate-limiting step is unimolecular, with A as the sole reactant.

           Hence, rate=k[A]=0.0030 M/s    where 'k' is the rate constant

(a) If [A] is doubled then rate=k[2A]=2k[A]=2*(0.0030 M/s)=0.0060M/s         Hence, rate will be doubled.

(b) As there is no [B] term in the rate expression, there will be no change in the rate even if [B] is doubled

(c) Both [A] and [B] are increased 5 times, Here, only the increase in [A] matters in the rate. [B] has no effect in the in the rate determining step.

     Hence, rate=k[5A]=5k[A]=5*(0.0030 M/s)=0.015M/s